OpenStudy (anonymous):

how do i verify (cos x/ 1 + sin x) + (1 + sin x/ cos x) = 2 sec x? i don't know how to verify anything at all can someone explain how it works?

Nnesha (nnesha):

\[\huge\rm \frac{ \cos(x) }{ 1+\sin(x) } +\frac{ 1+\sin(x) }{ \cos(x) }\] 1)what is the common denominator ?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

i really dont know anything about this

Nnesha (nnesha):

alright let's change that to `simple` algebra

Nnesha (nnesha):

\[\huge\rm \frac{ x }{ y }+\frac{ y }{ x }\] what's the common denominator ?

OpenStudy (anonymous):

xy

Nnesha (nnesha):

okay let's say y represent 1+sinx and x = cos(x) \[y=1+\sin(x)\] \[x=\cos(x)\] now can you tell me what would be the common denominator\[\huge\rm \frac{ \cos(x) }{ 1+\sin(x) } +\frac{ 1+\sin(x) }{ \cos(x) }\] for this question :=)

OpenStudy (anonymous):

1+sinx*cosx ?

Nnesha (nnesha):

PERFECT!

OpenStudy (anonymous):

hi i speak Arabic

OpenStudy (anonymous):

salam alakom

OpenStudy (anonymous):

Nnesha

Nnesha (nnesha):

\[\huge\rm \frac{ \cos(x) }{ \color{red}{1+\sin(x)} } +\frac{ 1+\sin(x) }{\color{blue}{ \cos(x)} }\] \[\huge\rm \frac{ \cos(\color{blue}{cos}) + (\color{reD}{1+\sin})(1+\sin)}{ \cos(x) (1+\sin(x))}\] multiply the numerator of `1st fraction` by the `denominator` of 2nd fraction multiply the numerator of `2nd fraction` with the denominator of `1st fraction`

OpenStudy (anonymous):

so would it become cos^2x and (1+sin)^2 on the numerator??

Nnesha (nnesha):

yes right!!!

Nnesha (nnesha):

\[(1+\sin(x))^2 = (1+\sin)(1+\sin)\] (1+sin)^2 is same as (1+sin)(1+sin) foil it

OpenStudy (anonymous):

and what happens when i foil it?

Nnesha (nnesha):

you will get 3 terms :3

Nnesha (nnesha):

then you will be able to cancel out some stuff at the numerator :=)

OpenStudy (anonymous):

so 1+2sinx+sin^2x ???

Nnesha (nnesha):

yes right \[\huge\rm \frac{ \color{ReD}{cos^2x}+1+2sinx+\sin^2x }{ cosx(1+sinx) }\] cos^2x= what ? familiar with the identity ?

Nnesha (nnesha):

\[\huge\rm sin^2 \theta+\cos^2\theta=1\] this one solve for cos^2

OpenStudy (anonymous):

cos^2=1-sin^2 ?

Nnesha (nnesha):

replace cos^2 with 1-sin^2 and then simplify !

Nnesha (nnesha):

\[\huge\rm \frac{ \color{ReD}{1-sin^2x}+1+2sinx+\sin^2x }{ cosx(1+sinx) }\]

OpenStudy (anonymous):

2+2sinx ?

Nnesha (nnesha):

yes! what is common factor ?

Nnesha (nnesha):

\[\huge\rm \frac{ 2+2sinx }{ \cos(x)(1+sinx) }\] 2+sinx what is common in both terms?

OpenStudy (anonymous):

sinx?

Nnesha (nnesha):

no first term is 2 and 2nd one is 2sinx so sinx isn't common

OpenStudy (anonymous):

so 2

Nnesha (nnesha):

yes right take out the 2 what will you get ?

OpenStudy (anonymous):

sinx

Nnesha (nnesha):

nah

Nnesha (nnesha):

when you take out 2 from 2 what will you have left wth?

OpenStudy (anonymous):

0

Nnesha (nnesha):

haha well no in other words you have to divide both terms with the common factor so \[\huge\rm \frac{ 2}{ 2} +\frac{ 2sinx }{ 2 } =2(??+???)\]

OpenStudy (anonymous):

1+sinx ?

Nnesha (nnesha):

yes! and we should keep the common factor outside the parentheses \[\huge\rm \frac{ 2(1+sinx) }{ cosx(1+sinx) }\]

Nnesha (nnesha):

simplify! that

OpenStudy (anonymous):

2/cosx

Nnesha (nnesha):

yes right and last step ?

OpenStudy (anonymous):

i dont know

Nnesha (nnesha):

thats fine! 2/cosx is same as \[\huge\rm 2 \times \frac{ 1 }{ cosx }\]

Nnesha (nnesha):

and 1/cosx= ??

OpenStudy (anonymous):

secx

Nnesha (nnesha):

so 2/cosx= ? :=)

OpenStudy (anonymous):

EXCITING

OpenStudy (anonymous):

i honestly have no idea how to do anything with identities so thank you

Nnesha (nnesha):

ikr! trig is FUN(well i hate graph)

Nnesha (nnesha):

my pleasure! :=) welcome to openstudy!

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @rami2021 salam alakom \(\color{blue}{\text{End of Quote}}\) Walikum salam and yeah arabic is my fvt language but all i know is *kafya halik ?? ;~;

Nnesha (nnesha):

sorry about that alejand :3

OpenStudy (anonymous):

no problem!

OpenStudy (anonymous):

i'm about to post another question, i have one left